Enter the equation of the following using line in slope-intercept form.
Two points on the line: (4, -2) and (-2, 4).

Answer by nyc_function(2741)

Let m = slope
m = 4 - (-2)/-2 - 4
m = 4 + 2/-6
m = 6/-6
m = -1
y - y1 = m(x - x1)
y - (-2) = -1(x - 4)
y + 2 = -x + 4
y = -x - 4 - 2
y = -x - 6
Done!